Let X be a random variable with the following cumulative distribution function (CDF): y<0 (a) Whats P(X < 2)? (b) Whats P(X > 2)? c)Whats P(0.5 X < 2.5)? (d) Whats P(X 1)? (e) Let q be a number such that F(0.6. Whats q?
Answer #1

a)from above :P(X<=2) =F(2) =(2+1)/4 =3/4

b)P(X>2) =1-P(X<=2)=1-3/4=1/4

c)P(0.5<X<=2.5)=F(2.5)-F(0.5)=(2.5+1)/4-(0.5/2)=7/8-1/4=5/8

d)P(X=1)=P(X<=1)-P(X<1)=(1+1)/4-(1/2)=1/2-1/2=0

e)

here as F(q)=(q+1)/4 =0.6

q=1.4